OpenStudy (anonymous):

What is the exact value of cos75* cos15*+ sin75* sin15* *=Degrees

OpenStudy (anonymous):

note that cos(a) = sin(90-A)

OpenStudy (anonymous):

the value is cheddar cheese

OpenStudy (anonymous):

so Cos 75=sin(90-75).......... i am being taught with the formulas sin (A+B)= sinA cosB + cosA sinB also cos (A+B)= cosA cosB - sinAsinB.... Do you know that one i know i have to use a 30 45 60 chart of sin cos and tan with raidiacles and franctions and fill in i just dont know how ...

OpenStudy (anonymous):

chedddar cheese

OpenStudy (anonymous):

Use the following identity to find the value: \[\cos(A-B)=\cos A \cos B+\sin A \sin B\]

OpenStudy (anonymous):

ah yes, anwar is right, ignore my answer. I spaced out on the identity stuff.

OpenStudy (anonymous):

^^ I am always right :P

OpenStudy (anonymous):

So, what do you think the value would be?

OpenStudy (anonymous):

Thank you for helping though still :] Hhaha ok now am i right about the chart thing..